OpenStudy (anonymous):

Part 1: Solve the inequality -3x + 4 <= x + 8 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1.

OpenStudy (anonymous):

x>= -1

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

intersection of both graphs is at (-1,7)

OpenStudy (anonymous):

explain plz

OpenStudy (anonymous):

since x+8 has positive slope u can say x>=1

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

\[-3x+4\leq x+8\] \[4\leq4x+8\] \[-4\leq 4x\] \[-1\leq x\]

OpenStudy (anonymous):

x is greater than or equal to -1